# Algebraic power series and their automatic complexity II: modulo prime powers

### v2

@@ Abstract.
Christol and, independently, Denef and Lipshitz showed that an algebraic sequence of $p$-adic integers (or integers) is $p$-automatic when reduced modulo $p^{\alpha}$. Previously, the best known bound on the minimal automaton size for such a sequence was doubly exponential in $\alpha$. We improve this bound to the order of $p^{\alpha^{3}hd}$, where $h$ and $d$ are the height and degree of the minimal annihilating polynomial modulo $p$. We achieve this bound by showing that all states in the automaton are naturally represented in a new numeration system. This significantly restricts the set of possible states. Since our approach embeds algebraic sequences as diagonals of rational functions, we also obtain bounds more generally for diagonals of multivariate rational functions.
@@ 1. Introduction
Christol [6] and, using a different approach, Denef and Lipshitz [8] proved a generalization of Christol’s theorem for power series with coefficients in $\mathbb{Z}$ or, more generally, the set $\mathbb{Z}_{p}$ of $p$-adic integers where $p$ is a prime. Namely, if $F=\sum_{n\geq 0}a(n)x^{n}\in\mathbb{Z}_{p}\llbracket x\rrbracket$ is algebraic, then the sequence $(a(n)\bmod p^{\alpha})_{n\geq 0}$ of elements in the ring $\mathcal{R}_{p^{\alpha}}\mathrel{\mathop{:}}=\mathbb{Z}/p^{\alpha}\mathbb{Z}$ is $p$-automatic for each $\alpha\geq 1$. In this article, we bound the size of the minimal automaton generating this sequence. Here and throughout this article, automata read the least significant digit first. From one of the constructions of Denef and Lipshitz [8, Remark 6.6], one obtains that the number of states is at most $p^{\alpha(p^{\alpha-1}\max(h,d)+1)^{2}}$ [17, Remark 2.2]. This bound is doubly exponential in $\alpha$. The techniques from [16] can be generalized to give a bound in this setting, but this bound is also doubly exponential. In this article, we significantly reduce the bound to roughly $p^{\alpha^{3}hd}$. This reduction is achieved primarily by identifying new structure in the representations of the states of the automaton. To state the main results, we introduce the following terminology.
@@
Let $P\in\mathbb{Z}_{p}[x,y]$ such that $P(0,0)=0$ and $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$. The Furstenberg series associated with $P$ is the unique element $F\in\mathbb{Z}_{p}\llbracket x\rrbracket$ satisfying $F(0)=0$ and $P(x,F)=0$.
The condition $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ says that the coefficient of $x^{0}y^{1}$ in $P$ is nonzero modulo $p$ and implies that $\deg_{y}(P\bmod p)\geq 1$. If $\deg_{x}P=0$, then $F$ is the trivial 00 series; we exclude it since otherwise the statement of the following theorem does not hold when $h=0$.
@@
Theorem 1 follows from a finer result, whose statement involves the following functions. Define $\operatorname{parts}(n)$ to be the set of all integer partitions of $n$. The Landau function $g(n)$ is the maximum value of $\operatorname{lcm}(\sigma)$ over all integer partitions $\sigma\in\operatorname{parts}(n)$ [20, A000793]. For example, $g(5)$ is the maximum value among $\operatorname{lcm}(5)$, $\operatorname{lcm}(4,1)$, $\operatorname{lcm}(3,2)$, $\operatorname{lcm}(3,1,1)$, $\operatorname{lcm}(2,2,1)$, $\operatorname{lcm}(2,1,1,1)$, and $\operatorname{lcm}(1,1,1,1,1)$, so $g(5)=6$. Finally, define
@@
$p^{N}+p^{N-\alpha(\alpha+1)(h+d-1)/2}\mathcal{L}(h,d,d)\\ +\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{t}})+\left\lceil\log_{p}\max(h,d-1)\right\rceil+\max(\alpha,2(\alpha-1))+\tfrac{p^{u}-1}{p-1}.$
We can recover the bound for a sequence of elements in the field $\mathbb{F}_{p}$ [16, Theorem 1] as follows. Let $\alpha=1$, and let $P\in\mathbb{F}_{p}[x,y]$ such that $\deg_{x}P=h$ and $\deg_{y}P=d$, so that $u=1$. Then the bound in Theorem 2 is at most
$p^{hd}+p^{(h-1)(d-1)}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}\max(h,d-1)\right\rfloor+1+\left\lceil\log_{p}\max(h,d-1)\right\rceil+1+1.$
@@
(1) $p^{hd}+p^{(h-1)(d-1)}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}h\right\rfloor+\left\lceil\log_{p}\max(h,d-1)\right\rceil+3$
in the previous article [16].
@@
There are other benefits of base-$\frac{p}{Q}$ representations as well. Representations of states in this numeration system are much more compact than their full expansions as Laurent polynomials, so using base-$\frac{p}{Q}$ representations greatly the reduces the amount of memory and time required to compute an automaton. For the sequence of Catalan numbers $C(n)_{n\geq 0}$, previously we were able to compute an automaton for $(C(n)\bmod 2^{9})_{n\geq 0}$, which has $2403$ states, but not for larger powers of $2$ [17]. With base-$\frac{p}{Q}$ representations, we are able to compute an automaton for $(C(n)\bmod 2^{14})_{n\geq 0}$; it has $174037$ states and required $34$GB of memory. From this automaton, one computes that only $\frac{2990}{2^{14}}\approx 18.2\%$ of the residues modulo $2^{14}$ are attained by the Catalan numbers, and only $\frac{2037}{2^{14}}\approx 12.4\%$ of residues are attained infinitely many times. This agrees with results of Straub [21, Table 2], who previously computed the residues attained by $C(n)$ modulo $2^{14}$ by representing $C(n)$ as a constant term and using scaling $p$-schemes rather than $p$-automata. In particular, these computations establish an upper bound of $0.124$ on the density of the set $\{C(n):n\geq 0\}$ in $\mathbb{Z}_{2}$. It is not known whether this density is positive.
@@
The second step is to obtain a preliminary upper bound on the automaton size. This is the subject of Section 4. We define a vector space $\mathcal{W}$ and show that $\mathcal{W}$ contains the base-$\frac{p}{Q}$ representations of most of the automaton states. To do this, we bound the degrees of the digits of the base-$\frac{p}{Q}$ representations of states. The space $\mathcal{W}$ has dimension $N=\frac{1}{6}\alpha(\alpha+1)((2hd-1)\alpha+hd+1)$, and this gives the main term in Theorem 2. The only states whose base-$\frac{p}{Q}$ representations do not belong to $\mathcal{W}$ are those reachable from the initial state by either a fixed small number of transitions or reading a sequence of 00s. The former are counted easily, and the latter correspond to the states in the orbit of the initial state under a certain linear transformation $\lambda_{0,0}$.
@@
Bounding the period lengths of coefficients of rational power series in Section 6 and the orbit size under $\lambda_{0}$ in Section 7 are considerably more involved over $\mathcal{R}_{p^{\alpha}}$ than in the proof of (1), which relied on facts about finite fields. One difficulty is that $\mathcal{R}_{p^{\alpha}}[z]$ does not have unique factorization. Fortunately, the first bound essentially depends on the period length of the coefficient sequence modulo $p$. Also, we use a lifting-the-exponent lemma throughout to show that the second bound essentially depends on information modulo $p$.
In Section 9, we generalize from algebraic series to diagonals of rational functions. The proof of Theorem 2 begins by converting the algebraic power series $F$ to the diagonal of a rational function in $2$ variables. By starting directly with the latter, the same approach allows us to bound the automaton size for the diagonal of a rational function in $m$ variables modulo $p^{\alpha}$, where the diagonal operator $\mathcal{D}$ is defined on multivariate series analogously to bivariate series.
@@
The dominant factor in the previous bound is $p^{\alpha^{m+1}h_{1}\cdots h_{m}}$. When $m=2$, we improve Theorem 3 to obtain the asymptotic bound $(1+o(1))p^{\alpha(\alpha+1)(2\alpha+1)hd/6}$ in Theorem 46. However, this approach does not currently extend to $m\geq 3$.
For polynomials $P,Q$ with integer coefficients, Beukers [3, Corollary 4.2] showed that if $\deg_{x_{i}}P\leq\deg_{x_{i}}Q=\mathrel{\mathop{:}}h_{i}$ and $\alpha-1\leq\rho-\left\lceil\rho/p\right\rceil$ for some integer $\rho$, then there is an automaton generating $(a(n)\bmod p^{\alpha})_{n\geq 0}$ with size at most $p^{\alpha(\rho h_{1}+1)\cdots(\rho h_{m}+1)}$. In particular, if $\alpha\leq p$ then we can take $\rho=\alpha$ and obtain the bound $p^{\alpha(\alpha h_{1}+1)\cdots(\alpha h_{m}+1)}$, which has the same dominant factor as the bound in Theorem 3 if $\deg_{x_{i}}Q=\deg_{x_{i}}(Q\bmod p)$ and $\deg_{x_{i}}P=\deg_{x_{i}}(P\bmod p)$, but which is asymptotically larger. Adamczewski, Bostan, and Caruso [1, Corollary 1.4] investigated the automatic complexity of multivariate algebraic series $F$ over a perfect field of positive characteristic. Taking the field to be $\mathbb{F}_{p}$ so that $\alpha=1$, they obtained the same dimension $M=\prod_{i=1}^{m}(h_{i}+1)$ as in Theorem 3, with a further refinement involving the total height of the annihilating polynomial of $F$.
Finally, in Section 10, we give another application of the base-$\frac{p}{Q}$ numeration system. For a given algebraic power series $F$, the automaton for $F\bmod p^{\alpha}$ projects naturally to the automaton for $F\bmod p^{\beta}$ for each $\beta\leq\alpha$, as was shown by the authors [18]. Consequently, the inverse limit as $\alpha\to\infty$ of these automata exists. However, no explicit description of its states was known. We show that the base-$\frac{p}{Q}$ representations of the states of the automaton for $F\bmod p^{\beta}$ are simply truncations of those for $F\bmod p^{\alpha}$. This gives explicit descriptions of the states of the inverse limit automaton as inverse limits of finite base-$\frac{p}{Q}$ representations.
@@ 2. The module of possible states
Theorems 1 and 2 are concerned with automatic sequences with elements in $\mathcal{R}_{p^{\alpha}}$. These sequences arise in the following result of Denef and Lipshitz [8, Lemma 6.3], which extends Furstenberg’s theorem [10] to an integral domain. We state it for the $p$-adic integers, along with an automaticity result proved by Denef and Lipshitz [8, Remark 6.6]; see also [17, Theorem 2.1]. The diagonal operator $\mathcal{D}$, acting on bivariate power series, is defined by
@@
The smallest automaton that generates $a(n)_{n\geq 0}$ and that is not affected by leading 00s is its minimal automaton. Eilenberg’s theorem gives a bijection between the states of the minimal automaton and the elements of the $p$-kernel.
@@
A proof of Proposition 5 can be found in [17, Proposition 1.9]. It uses the following lifting-the-exponent lemma for (Laurent) polynomials, whose proof is analogous to the proof for integers. We will use Lemma 6 throughout this article.
@@
If $R,\bar{R}\in\mathbb{Z}_{p}[x,x^{-1}]$ and $R\equiv\bar{R}\mod p$, then $R^{p^{\alpha-1}}\equiv\bar{R}^{p^{\alpha-1}}\mod p^{\alpha}$. Similarly, if $Q,\bar{Q}\in\mathbb{Z}_{p}[x,x^{-1},y,y^{-1}]$ and $Q\equiv\bar{Q}\mod p$, then $Q^{p^{\alpha-1}}\equiv\bar{Q}^{p^{\alpha-1}}\mod p^{\alpha}$.
The final step is to establish a connection between a power series $\sum_{n\geq 0}a(p^{e}n+r)x^{n}$ corresponding to a kernel sequence and a Laurent polynomial. This Laurent polynomial will be the numerator of a rational function whose diagonal is the desired power series. We do this using Theorem 4. We shear bivariate series by replacing $x$ with $xy^{-1}$. When we do this, the diagonal operator is replaced by the center row operator $\mathcal{C}$, defined by
@@
The condition $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ implies $d\geq 1$. Furthermore, if $F\bmod p^{\alpha}$ is not a polynomial then $h\geq 1$. This is because $h=0$ implies that the coefficient of each monomial $x^{i}y^{j}$ in $P$ with $i\geq 1$ is 00 modulo $p$. Therefore the denominator $P(xy,y)/y$ in Theorem 4 also has this property. Using the geometric series formula to expand the rational expression, we obtain a bivariate series that contains only finitely many nonzero diagonal monomials modulo $p^{\alpha}$.
@@
The constructed automaton is not sensitive to leading 00s.
@@
Let $S\in\mathcal{M}_{p^{\alpha}}$, let $x^{i}y^{j}$ be a stripped monomial in $S$ (that is, a monomial without its coefficient), and let $x^{I}y^{J}$ be a stripped monomial in $Q^{p^{\alpha}-p^{\alpha-1}}$; we are interested in products $x^{i}y^{j}\cdot x^{I}y^{J}$ which equal $1$, since these contribute to the constant term and therefore the output value. This implies $I=-i$ and $J=-j$. Since $i\geq 0$ and $I\geq 0$, we have $i=0=-I$. We consider the three cases $j<0$, $j>0$, and $j=0$. If $j>0$, then $J<0$; this implies $I>0$ by the condition $P(0,0)\neq 0$, which contradicts $I=0$. If $j<0$, then $-i\leq j<0$ by Lemma 9, which implies $i<0$ and contradicts $i=0$. Therefore $j=0=J$, so the constant term of $SQ^{p^{\alpha}-p^{\alpha-1}}$ is the product of the constant term of $S$ and the constant term of $Q^{p^{\alpha}-p^{\alpha-1}}$. The assumption $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ implies that the constant term $c\,x^{0}y^{0}$ of $Q$ is nonzero modulo $p$. It follows that the constant term of $Q^{p^{\alpha}-p^{\alpha-1}}$ is $c^{p^{\alpha}-p^{\alpha-1}}=1$, since the Euler totient function satisfies $\phi(p^{\alpha})=p^{\alpha}-p^{\alpha-1}$. Therefore the constant term of $\lambda_{0,0}(S)$ divided by the constant term of $Q$, which is the output assigned to the state $\lambda_{0,0}(S)$, is the constant term of $S$ divided by the constant term of $Q$. ∎

### v3

# Algebraic power series and their automatic complexity modulo prime powers

@@ Abstract.
Christol and, independently, Denef and Lipshitz showed that an algebraic sequence of $p$-adic integers (or integers) is $p$-automatic when reduced modulo $p^{\alpha}$. Previously, the best known bound on the minimal automaton size for such a sequence was doubly exponential in $\alpha$. Under mild conditions, we improve this bound to the order of $p^{\alpha^{3}hd}$, where $h$ and $d$ are the height and degree of the minimal annihilating polynomial modulo $p$. We achieve this bound by showing that all states in the automaton are naturally represented in a new numeration system. This significantly restricts the set of possible states. Since our approach embeds algebraic sequences as diagonals of rational functions, we also obtain bounds more generally for diagonals of multivariate rational functions.
@@ 1. Introduction
Christol [8] and, using a different approach, Denef and Lipshitz [10] proved a generalization of Christol’s theorem for power series with coefficients in $\mathbb{Z}$ or, more generally, the set $\mathbb{Z}_{p}$ of $p$-adic integers where $p$ is a prime. Namely, if $F=\sum_{n\geq 0}a(n)x^{n}\in\mathbb{Z}_{p}\llbracket x\rrbracket$ is algebraic, then the sequence $(a(n)\bmod p^{\alpha})_{n\geq 0}$ of elements in the ring $\mathcal{R}_{p^{\alpha}}\mathrel{\mathop{:}}=\mathbb{Z}/p^{\alpha}\mathbb{Z}$ is $p$-automatic for each $\alpha\geq 1$.
In this article, we bound the size of the minimal automaton generating this sequence. Here and throughout this article, automata read the least significant digit first. Our main results concern series defined as follows.
@@
Let $P\in\mathbb{Z}_{p}[x,y]$ such that $P(0,0)=0$ and $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$, that is, $\frac{\partial P}{\partial y}(0,0)$ is a unit. The Furstenberg series associated with $P$ is the unique element $F\in\mathbb{Z}_{p}\llbracket x\rrbracket$ satisfying $F(0)=0$ and $P(x,F)=0$.
From one of the constructions of Denef and Lipshitz [10, Remark 6.6], for Furstenberg series one obtains that the number of states is at most $p^{\alpha(p^{\alpha-1}\max(h,d)+1)^{2}}$ [19, Remark 2.2]. This bound is doubly exponential in $\alpha$. The techniques from [18] can be generalized to give a bound in this setting, but this bound is also doubly exponential. In this article, we significantly reduce the bound to roughly $p^{\alpha^{3}hd}$. This reduction is achieved primarily by identifying new structure in the representations of the states of the automaton.
In the definition of a Furstenberg series, the condition $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ says that the coefficient of $x^{0}y^{1}$ in $P$ is nonzero modulo $p$, and it implies that $\deg_{y}(P\bmod p)\geq 1$. If $\deg_{x}P=0$, then $F$ is the trivial 0 series; we exclude it since otherwise the statement of the following theorem does not hold when $h=0$.
@@
Theorem 1 follows from a finer result, whose statement involves the following functions. Define $\operatorname{parts}(n)$ to be the set of all integer partitions of $n$. The Landau function $g(n)$ is the maximum value of $\operatorname{lcm}(\sigma)$ over all integer partitions $\sigma\in\operatorname{parts}(n)$ [22, A000793]. For example, $g(5)$ is the maximum value among $\operatorname{lcm}(5)$, $\operatorname{lcm}(4,1)$, $\operatorname{lcm}(3,2)$, $\operatorname{lcm}(3,1,1)$, $\operatorname{lcm}(2,2,1)$, $\operatorname{lcm}(2,1,1,1)$, and $\operatorname{lcm}(1,1,1,1,1)$, so $g(5)=6$. Finally, define
@@
$p^{N}+p^{N-\alpha(\alpha+1)(h+d-1)/2}\mathcal{L}(h,d,d)\\ +\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{t}})+\left\lfloor\log_{p}\max(h,d)\right\rfloor+2\alpha-1+\tfrac{p^{u}-1}{p-1}.$
For a random assignment of the coefficients of $P\bmod p$, the value of $\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{t}})$ is typically $1$ when $p$ is large. For all polynomials $P$, using $\max(u_{\ell},u_{\textnormal{r}},u_{\textnormal{t}})\leq\left\lfloor\log_{p}\max(h,d)\right\rfloor+\alpha$, we obtain the simpler bound
$p^{N}+p^{N-\alpha(\alpha+1)(h+d-1)/2}\mathcal{L}(h,d,d)+2\left\lfloor\log_{p}\max(h,d)\right\rfloor+3\alpha-1+\tfrac{p^{u}-1}{p-1}.$
We recover the bound for a sequence of elements in the field $\mathbb{F}_{p}$ [18, Theorem 1] as follows. Let $\alpha=1$, and let $P\in\mathbb{F}_{p}[x,y]$ such that $\deg_{x}P=h$ and $\deg_{y}P=d$, so that $u=1$. Then the bound in Theorem 2 is at most
$p^{hd}+p^{(h-1)(d-1)}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}\max(h,d-1)\right\rfloor+1+\left\lfloor\log_{p}\max(h,d)\right\rfloor+1+1.$
@@
(1) $p^{hd}+p^{(h-1)(d-1)}\mathcal{L}(h,d,d)+\left\lfloor\log_{p}h\right\rfloor+\left\lfloor\log_{p}\max(h,d)\right\rfloor+3$
in [18].
@@
There are other benefits of base-$\frac{p}{Q}$ representations as well. Representations of states in this numeration system are much more compact than their full expansions as Laurent polynomials, so using base-$\frac{p}{Q}$ representations greatly reduces the amount of memory and time required to compute an automaton. For the sequence of Catalan numbers $C(n)_{n\geq 0}$, previously we were able to compute an automaton for $(C(n)\bmod 2^{9})_{n\geq 0}$, which has $2403$ states, but not for larger powers of $2$ [19]. With base-$\frac{p}{Q}$ representations, we are able to compute an automaton for $(C(n)\bmod 2^{14})_{n\geq 0}$; it has $174037$ states and required 3434GB of memory. From this automaton, one computes that only $\frac{2990}{2^{14}}\approx 18.2\%$ of the residues modulo $2^{14}$ are attained by the Catalan numbers, and only $\frac{2037}{2^{14}}\approx 12.4\%$ of residues are attained infinitely many times. This agrees with results of Straub [23, Table 2], who previously computed the residues attained by $C(n)$ modulo $2^{14}$ by representing $C(n)$ as a constant term and using scaling $p$-schemes rather than $p$-automata. In particular, these computations establish an upper bound of $0.124$ on the density of the set $\{C(n):n\geq 0\}$ in $\mathbb{Z}_{2}$. It is not known whether this density is positive.
Unrelated to our main results, in Section 4 we give another application of the base-$\frac{p}{Q}$ numeration system. For a given algebraic power series $F$, the automaton for $F\bmod p^{\alpha}$ projects naturally to the automaton for $F\bmod p^{\beta}$ for each $\beta\leq\alpha$, as was shown by the authors [20]. Consequently, the inverse limit as $\alpha\to\infty$ of these automata exists. However, no explicit description of its states was known. We show that the base-$\frac{p}{Q}$ representations of the states of the automaton for $F\bmod p^{\beta}$ are simply truncations of those for $F\bmod p^{\alpha}$. This gives explicit descriptions of the states of the inverse limit automaton as inverse limits of finite base-$\frac{p}{Q}$ representations.
Theorem 1 requires $F$, annihilated by $P$, to be a Furstenberg series. In Section 10, we consider the possibility that $F$ is not a Furstenberg series. A standard modification of $F$ gives an algebraic series $G$ whose annihilating polynomial has a nonzero derivative at the origin. If, in addition, this derivative is nonzero modulo $p$, then we obtain an asymptotic bound in Theorem 51; this bound is on the order of $p^{\alpha^{3}hd^{3}}$ where $h=\deg_{x}P$ and $d=\deg_{y}P$, so it is slightly larger than the bound in Theorem 1. However, the relevant derivative may be zero modulo $p$. For example, the polynomial $P=(x+1)(3x-1)y^{2}+1$ annihilates the generating series of the central trinomial coefficients, and for $p=2$ the technique does not work [19, Section 4.2].
@@
The second step is to obtain a preliminary upper bound on the automaton size. This is the subject of Section 5. We define a vector space $\mathcal{W}$ and show that $\mathcal{W}$ contains the base-$\frac{p}{Q}$ representations of most of the automaton states. To do this, we bound the degrees of the digits of the base-$\frac{p}{Q}$ representations of states. The space $\mathcal{W}$ has dimension $N=\frac{1}{6}\alpha(\alpha+1)((2hd-1)\alpha+hd+1)$, and this gives the main term in Theorem 2. The only states whose base-$\frac{p}{Q}$ representations do not belong to $\mathcal{W}$ are those reachable from the initial state by either a fixed small number of transitions or reading a sequence of 0s. The former are counted easily, and the latter correspond to the states in the orbit of the initial state under a certain linear transformation $\lambda_{0,0}$.
@@
Bounding the period lengths of coefficients of rational power series in Section 7 and the orbit size under $\lambda_{0}$ in Section 8 are considerably more involved over $\mathcal{R}_{p^{\alpha}}$ than in the proof of the bound (1), which relied on facts about finite fields. One difficulty is that $\mathcal{R}_{p^{\alpha}}[z]$ does not have unique factorization. We get around this by showing that the first bound essentially depends on the period length of the coefficient sequence modulo $p$. We use a lifting-the-exponent lemma throughout to show that the second bound also essentially depends on information modulo $p$.
Finally, in Section 11, we generalize Theorem 2 from algebraic series to diagonals of rational functions in several variables. The proof of Theorem 2 begins by converting the algebraic power series $F$ to the diagonal of a rational function (in $2$ variables). By starting directly with the latter, the same approach allows us to bound the automaton size for the diagonal of a rational function in $m$ variables modulo $p^{\alpha}$, where the diagonal operator $\mathcal{D}$ is defined on multivariate series analogously to bivariate series.
@@
The dominant factor in the previous bound is $p^{\alpha^{m+1}h_{1}\cdots h_{m}}$. For polynomials $P,Q$ with integer coefficients, a result of Beukers [5, Corollary 4.2] implies that if $\deg_{x_{i}}P\leq\deg_{x_{i}}Q=\mathrel{\mathop{:}}h_{i}$ and $\alpha-1\leq\rho-\left\lceil\rho/p\right\rceil$ for some integer $\rho$ then there is an automaton generating $(a(n)\bmod p^{\alpha})_{n\geq 0}$ with size at most $p^{\alpha(\rho h_{1}+1)\cdots(\rho h_{m}+1)}$. In particular, if $\alpha\leq p$ then we can take $\rho=\alpha$ and obtain the bound $p^{\alpha(\alpha h_{1}+1)\cdots(\alpha h_{m}+1)}$, which has the same dominant factor as the bound in Theorem 3 if $\deg_{x_{i}}Q=\deg_{x_{i}}(Q\bmod p)$ and $\deg_{x_{i}}P=\deg_{x_{i}}(P\bmod p)$, but which is asymptotically larger. Adamczewski, Bostan, and Caruso [2, Corollary 1.4] investigated the automatic complexity of multivariate algebraic series $F$ over a perfect field of positive characteristic. Taking the field to be $\mathbb{F}_{p}$ so that $\alpha=1$, they obtained the same dimension $M=\prod_{i=1}^{m}(h_{i}+1)$ as in Theorem 3, with a further refinement involving the total height of the annihilating polynomial of $F$.
For $m=2$, in Theorem 56 we improve Theorem 3 to obtain the asymptotic bound $(1+o(1))p^{\alpha(\alpha+1)(2\alpha+1)hd/6}$. New techniques would be required to extend this approach to $m\geq 3$.
@@ 2. The module of possible states
Theorems 1 and 2 are concerned with automatic sequences with elements in $\mathcal{R}_{p^{\alpha}}$. These sequences arise in the following result of Denef and Lipshitz [10, Lemma 6.3], which extends Furstenberg’s theorem [12] to an integral domain. We state it for the $p$-adic integers, along with an automaticity result proved by Denef and Lipshitz [10, Remark 6.6]; see also [19, Theorem 2.1]. The diagonal operator $\mathcal{D}$, acting on bivariate power series, is defined by
@@
The smallest automaton that generates $a(n)_{n\geq 0}$ and that is not affected by leading 0s is its minimal automaton. Eilenberg’s theorem gives a bijection between the states of the minimal automaton and the elements of the $p$-kernel.
@@
A proof of Proposition 5 can be found in [19, Proposition 1.9]. It uses the following lifting-the-exponent lemma for (Laurent) polynomials, whose proof is analogous to the proof for integers. We will use Lemma 6 throughout this article.
@@
If $Q,\bar{Q}\in\mathcal{R}_{p^{\alpha}}[x_{1},x_{1}^{-1},\dots,x_{m},x_{m}^{-1}]$ and $Q\equiv\bar{Q}\mod p$, then $Q^{p^{\alpha-1}}=\bar{Q}^{p^{\alpha-1}}$.
The final step is to establish a connection between a power series $\sum_{n\geq 0}a(p^{e}n+r)x^{n}$ corresponding to a kernel sequence and a (not necessarily unique) Laurent polynomial. This Laurent polynomial will be the numerator of a rational function whose diagonal is the desired power series. We do this using Theorem 4. We shear bivariate series by replacing $x$ with $xy^{-1}$. When we do this, the diagonal operator is replaced by the center row operator $\mathcal{C}$, defined by
@@
The condition $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ implies $d\geq 1$. Furthermore, if $F\bmod p^{\alpha}$ is not a polynomial then $h\geq 1$. This is because $h=0$ implies that the coefficient of each monomial $x^{i}y^{j}$ in $P$ with $i\geq 1$ is 0 modulo $p$. Therefore the denominator $P(xy,y)/y$ in Theorem 4 also has this property. Using the geometric series formula to expand the rational expression, we obtain a bivariate series that contains only finitely many nonzero diagonal monomials modulo $p^{\alpha}$.
@@
The constructed automaton is not sensitive to leading 0s.
@@
Let $S\in\mathcal{M}_{p^{\alpha}}$, let $x^{i}y^{j}$ be a stripped monomial in $S$ (that is, a monomial without its coefficient), and let $x^{I}y^{J}$ be a stripped monomial in $Q^{p^{\alpha}-p^{\alpha-1}}$; we are interested in products $x^{i}y^{j}\cdot x^{I}y^{J}$ which equal $1$, since these contribute to the constant term and therefore the output value. This implies I=iI=-i and J=jJ=-j. Since $i\geq 0$ and $I\geq 0$, we have i=0=Ii=0=-I. We consider the three cases $j<0$, $j>0$, and $j=0$. If $j>0$, then $J<0$; this implies $I>0$ by the condition $P(0,0)\neq 0$, which contradicts $I=0$. If $j<0$, then $i\leq j<0$ by Lemma 9, which implies $i<0$ and contradicts $i=0$. Therefore $j=0=J$, so the constant term of $SQ^{p^{\alpha}-p^{\alpha-1}}$ is the product of the constant term of $S$ and the constant term of $Q^{p^{\alpha}-p^{\alpha-1}}$. The assumption $\frac{\partial P}{\partial y}(0,0)\mathrel{\not\equiv}0\mod p$ implies that the constant term $c\,x^{0}y^{0}$ of $Q$ is nonzero modulo $p$. It follows that the constant term of $Q^{p^{\alpha}-p^{\alpha-1}}$ is $c^{p^{\alpha}-p^{\alpha-1}}=1$, since the Euler totient function satisfies $\phi(p^{\alpha})=p^{\alpha}-p^{\alpha-1}$. Therefore the constant term of $\lambda_{0,0}(S)$ divided by the constant term of $Q$, which is the output assigned to the state $\lambda_{0,0}(S)$, is the constant term of $S$ divided by the constant term of $Q$. ∎